OpenStudy (anonymous):

find k so that the following function is continuous on any interval: h(x) = kcosx on 0

OpenStudy (aum):

In order for h(x) to be continuous at \(x = 5\pi\),\[ \lim_{x \rightarrow 5\pi^-}h(x) = \lim_{x \rightarrow 5\pi^+}h(x) \\ \lim_{x \rightarrow 5\pi^-}k\cos(x) = \lim_{x \rightarrow 5\pi^+}(14-x) \\ k\cos(5\pi) = 14 - 5\pi \\ \cos(5\pi) = \cos(3\pi) = \cos(\pi) = -1 \\ -k = 14 - 5\pi \\ \\ k = 5\pi - 14 \]